OpenStudy (anonymous):

Find the area of the region enclosed by the functions y=x^6 y=x

OpenStudy (anonymous):

Easy first you need to find the two areas of inersection

OpenStudy (anonymous):

intersection*

OpenStudy (anonymous):

Then you integrate by those bounding areas and subtract the two functions

OpenStudy (anonymous):

is the area of intersection at 1,1??

OpenStudy (lgbasallote):

have you tried graphing this?

OpenStudy (anonymous):

yes i just did

OpenStudy (lgbasallote):

you can find the intersection by making them equal to one another \[x = x^6\] \[0 = x^6 - x\] \[0 = x(x^5 - 1)\] the intersection would be x = 0 and x = 1

OpenStudy (lgbasallote):

so what is the value of y when x = 0 y = x y = 0 when x = 1 y = x y = 1 therefore the intersections are (0,0) and (1,1)

OpenStudy (anonymous):

ok so ther is two intersections then

OpenStudy (anonymous):

but we use x as the bounds because the y are going to vary because of the functions

OpenStudy (lgbasallote):

correct me if im wrong but i think the y = x^6 graph is to the right of y=x right

OpenStudy (lgbasallote):

sp it would be \[\huge \int_0^1 (x^6 - x)dx\]

OpenStudy (lgbasallote):

so simply integrate/ \[\int_0^1 x^6dx - \int_0^1 xdx\]

OpenStudy (anonymous):

yes in the right

OpenStudy (anonymous):

No I think it's the other way around x-x^6

OpenStudy (lgbasallote):

do you get what im doing?

OpenStudy (anonymous):

yes i do

OpenStudy (lgbasallote):

but x^6 is on the right...right @Romero ?

OpenStudy (anonymous):

careful here

OpenStudy (lgbasallote):

i dont have a graph so it will help if there is one lol

OpenStudy (anonymous):

I see it as x^6 is under x

OpenStudy (anonymous):

on the interval \([0,1]\) we have \(x^6<x\)

OpenStudy (anonymous):

yes on the right

OpenStudy (anonymous):

you can imagine with out a graph what is bigger, \(\frac{1}{2}\) or \((\frac{1}{2})^6\)?

OpenStudy (anonymous):

1/2

OpenStudy (lgbasallote):

so i guess it's \[\int_0^1 (x-x^6)dx\] huh

OpenStudy (lgbasallote):

ok i admit i was wrong lol

OpenStudy (anonymous):

yeah sure and heck if you are going to ask wolfram just ask it is very accommodating http://www.wolframalpha.com/input/?i=area+between+y%3Dx^6+%2Cy%3Dx wont help on an exam though

OpenStudy (anonymous):

so it is not on the right

OpenStudy (lgbasallote):

@blopez y=x^6 is to the left

OpenStudy (lgbasallote):

i just checked the graph haha

OpenStudy (anonymous):

i would use the work "below"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Don't ask wolfram for answers... I really don't trust him -.-

OpenStudy (anonymous):

thanks for the help